### v1

# Long-Tailed Continual Learning For Visual Food Recognition

Abstract
Deep learning based food recognition has achieved remarkable progress in predicting food types given an eating occasion image. However, there are two major obstacles that hinder deployment in real world scenario. First, as new foods appear sequentially overtime, a trained model needs to learn the new classes continuously without causing catastrophic forgetting for already learned knowledge of existing food types. Second, the distribution of food images in real life is usually long-tailed as a small number of popular food types are consumed more frequently than others, which can vary in different populations. This requires the food recognition method to learn from class-imbalanced data by improving the generalization ability on instance-rare food classes. In this work, we focus on long-tailed continual learning and aim to address both aforementioned challenges. As existing long-tailed food image datasets only consider healthy people population, we introduce two new benchmark food image datasets, VFN-INSULIN and VFN-T2D, which exhibits on the real world food consumption for insulin takers and individuals with type 2 diabetes without taking insulin, respectively. We propose a novel end-to-end framework for long-tailed continual learning, which effectively addresses the catastrophic forgetting by applying an additional predictor for knowledge distillation to avoid misalignment of representation during continual learning. We also introduce a novel data augmentation technique by integrating class-activation-map (CAM) and CutMix, which significantly improves the generalization ability for instance-rare food classes to address the class-imbalance issue. The proposed method is evaluated on Food101-LT, VFN-LT, VFN-INSULIN and VFN-T2D and show promising performance with large margin improvements compared with existing methods. We conduct an ablation study and discuss potential techniques that can further improve the performance, demonstrating great potential to deploy our method in real world food recognition applications.
Index Terms:
Continual learning, long-tailed distribution, food recognition, knowledge distillation, data augmentation
I Introduction
The emergence of modern deep learning technologies have enabled the automatic food nutrition content analysis, including image-based dietary assessment [1, 2, 3, 4], to help monitor and inform interventions to improve dietary intake and to prevent chronic diseases such as diabetes. As the first and fundamental step of image-based dietary assessment, food recognition aims to identify food types given an input image and the overall dietary assessment performance greatly relies on the precise food recognition results. Though existing deep learning based food recognition methods [5, 6, 7, 8, 9, 10] have achieved remarkable performance by training off-the-shelf Convolutional Neural Networks (e.g. ResNet[11]) using static datasets (e.g. Food-101[12], Food2K [13]), there are still two major challenges when applying them in a real world scenario including (i) how to update the model when new food classes appear sequentially overtime, and (ii) how to address the severe class-imbalance issue since the real life food images are usually in long-tailed distribution as shown in [14] where a minority of foods classes (i.e. instance-rich or head classes) are consumed more frequently than the remaining majority food classes (i.e. instance-rare or tail classes). The food recognition performance could drop dramatically without considering both two challenges. As shown in Figure 1, an ideal food recognition system in the real world should be able to learn new foods incrementally in long-tailed distribution with class-imbalance training samples at each incremental phase.
Figure 1: The overview of long-tailed continual learning. New classes with imbalanced distribution arrive sequentially over time at each incremental learning phase. The updated model should be able to learn new classes continuously and perform classification on all the classes seen so far.
Continual learning, also known as incremental learning or lifelong learning, has been studied with the objective of learning new classes continuously without catastrophic forgetting [15] of learned knowledge. Compared with the naive solution of retraining the model from scratch whenever encountering a new class, continual learning is more practical as it only requires the data from new classes during the learning process, leading to the improvements in time, computation and memory efficiency [16]. However, the problem becomes more challenging if the data exhibits long-tailed distribution where the model needs to address both catastrophic forgetting and class-imbalance issues. Though the most recent work [17] aimed to solve this problem by introducing a 2-stage framework, the detached training process also presents challenges when implemented in real-world applications due to the inefficiency caused by manual fine-tuning of each individual stage. In addition, none of the existing work target on food images, which can be more challenging due to the intra-class variation and inter-class similarity.
Existing continual learning methods show the effectiveness of applying knowledge distillation and storing a small fixed number of seen images as exemplars to mitigate catastrophic forgetting. However, both techniques become less effective in long-tailed distribution. Specifically, the knowledge distillation [18] may even harm the performance when teacher’s model is not trained on balanced data due to the bias in output logits as shown in a recent study [19]. On the other hand, distilling knowledge through learned representations impose a new challenge of feature space misalignment [20] as the learned representation needs to evolve during continual learning to accommodate new classes. Regarding using an exemplar set, most classes in long-tailed distribution may contain only a few training samples. Consequently, the overall performance may still be hindered even when all available samples are stored for instance-rare classes due to the poor generalization ability.
In this work, we focus on designing an end-to-end long-tailed continual learning framework for visual food recognition. We leverage feature-based knowledge distillation while incorporating an additional prediction head that projects the current representation space to the past, which addresses the misalignment issue by providing more freedom to the student model and encourages the retention of the learned knowledge. In addition, inspired by the most recent work [21] that uses the context-rich information in head classes to help the tail classes, we introduce a new data augmentation technique by integrating class-activation-map (CAM) and CutMix [22], which cuts the most important region calculated by CAM in instance-rare classes data as foreground and pastes into the instance-rich classes images. With minimal computational overhead, this method significantly enhances the generalization capabilities of tail classes. We evaluate our method on existing long-tailed food image datasets including Food101-LT and VFN-LT. We further expand the VFN-LT to include two additional population groups, namely Insulin Takers and those with Type 2 diabetes without taking insulin, denoted as VFN-INSULIN and VFN-T2D, respectively. Our proposed framework achieves the best performance with a large improvement margin compared to existing methods while not requiring detached training stages. Finally, we conduct an ablation study to evaluate the effectiveness of each component in our proposed method and discuss potential techniques that can boost the accuracy for implementing in real world applications. The main contributions of this work are summarized as in the following:
  • We study the long-tailed continual learning in this work, which is closely related to food recognition in various real world scenarios. Despite its significance, this area of research is currently under-explored.
  • Two new benchmark long-tailed food image datasets including VFN-INSULIN and VFN-T2D are introduced, which exhibit the real world food consumption for the population of insulin takers and type 2 diabetes without taking insulin, respectively.
  • We propose a novel framework to address both catastrophic forgetting and class-imbalance effectively by using feature-based knowledge distillation with a predication head and a novel CAM-based CutMix for data augmentation. In addition, an integrated loss is introduced to strike the balance between learning new classes and maintaining the learned knowledge.
  • We conduct a set of comprehensive experiments on all long-tailed continual learning benchmarks for food recognition and discuss potential techniques that could boost the accuracy for facilitating the deployment in real world food recognition.
II Related Word
In this section, we summarize existing methods that are most related to our work including the food recognition, long-tailed recognition and continual learning.
II-A Food Recognition
Food image recognition is a challenging yet practical task that is related to various real world applications such as the image-based dietary assessment [23, 24, 25]. The performance of nutritional content analysis such as energy [26] or macronutrients are heavily reliant on the accuracy of recognizing and predicting the correct food types from food images. Most existing deep learning based work leverage off-the-shelf models [27, 28, 11, 29] and train on static food image datasets [12, 13, 30, 31, 32, 9]. In order to address the issue of inter-class similarity and intra-class variability, a manually built hierarchy based food recognition is proposed in [10]. Later, Mao et al[9] propose to construct a food hierarchy based on visual similarity and nutrition content [33] without requiring human efforts. In addition, food recognition has been studied under different scenarios such as ingredient recognition [34, 35], fine-grained recognition [36, 37, 38], few-shot learning [39, 40], long-tailed recognition [14, 41] and continual learning [42, 43]. However, none of the existing method are capable of learning new classes continuously in long-tailed distributions, which closely relates to real-world food recognition as new foods appear sequentially overtime with a minority of foods consumed more frequently than the others [14]. Though the most recent work [17] aims to integrate continual learning and long-tailed recognition, they decouple the training process to multi-stages and do not focus on food images. In this work, we target on long-tailed continual learning for visual food recognition to fill this gap and introduce a novel end-to-end framework to address both class-imbalance issue and catastrophic forgetting simultaneously.
II-B Long-tailed Recognition
Image recognition in long-tailed distribution has been widely studied in recent years [44]. Existing work can be mainly categorized into two main groups including re-weighting and re-sampling based methods. The major challenge of long-tailed recognition is the imbalance training data between instance-rich (head) and instance-rare (tail) classes. The re-weighting based methods aim to balance the loss or gradients during the training process. The class-level re-weighting loss assigns weights to each class such as the Balanced Softmax loss [45], Label-Distribution-Aware Margin loss [46] where the inverse of class frequency [47] is widely used as the class weights. In addition, re-sampling based techniques aim to construct a balanced training set by over-sampling the tail classes or under-sampling the head classes. However, the naive over-sampling [48] by simply repeating the tail classes data for balanced training further intensifies the over-fitting issue and the naive under-sampling [49] by randomly removing the training data of head classes results in performance degradation. Therefore, most existing work performs efficient data augmentation techniques to improve the generalization ability of tail classes and achieve better overall recognition performance. Gao et al[41] propose Dynamic Mixup for multi-label long-tailed recognition problem, which dynamically adjusts the selection of images based on the previous training performance and set the label of synthetic image as the union of two images. The most recent work proposed CMO [21], which applies CutMix [22] for data augmentation by cutting the foreground region in tail classes images and paste in head classes background. The center idea of CMO is to leverage the context rich information from the head classes to help the generalization of tail classes. Later, He et al[14] improves the CMO to use visually similar image pairs and allows for multi-image CutMix to achieve improved performance. In spite of the efficiency of CutMix for data augmentation, one of the limitations is that it suffers from loss of semantic information of the original image since the cut region is generated randomly. Inspired by [50], we introduce a novel CAM based CutMix, which is able to seamlessly combine the images without losing semantic information, which will be illustrated in Section IV.
II-C Continual Learning
Continual learning, also known as incremental learning or lifelong learning, has been studied under different scenarios such as class-incremental, task-incremental and domain-incremental learning as categorized in [51]. From the perspective of real world application, we primarily focus on class-incremental learning in this work, which aims to learn new classes continuously and perform classification on all classes seen so far during the inference phase, which does not require the task index or use multi-head classifier [52] compared to task-incremental learning. The domain-incremental learning instead does not involve new classes but only the domain shift for learned knowledge. The major challenge of class-incremental learning is catastrophic forgetting [15] where the updated model quickly forgets the learned knowledge due to the unavailability of learned classes data. There are two main groups of existing class-incremental learning methods including regularization based and replay based methods.
Regularization based methods address forgetting by restricting the change of learned parameters during the learning of new classes. The initial work freeze the learned parameters in fully connected layer [53] or constraining the weights update [54], which also confines the model’s ability to learn from new data. Later, Li et al[55] proposed to use knowledge distillation [18] to maintain the learned knowledge by mimicking the output logits distribution from a teacher model, which is then improved in [56] by adding stronger constraints. In addition, feature-based knowledge distillation for class-incremental learning are applied in [57] by minimizing the discrepancy of learned representations between student and teacher models, which is further developed in [58] integrating the logits and feature-based distillation. However, existing knowledge distillation methods are developed on balanced data and becomes less efficient in long-tailed distribution. As revealed in [19], the knowledge distillation using output logits may even harm the overall performance if the teacher model is not trained on balanced data due to the bias towards instance-rich classes. Furthermore, directly applying feature based distillation may also impose new challenge such as feature space misalignment [20] due to the evolving of feature space when learning new classes especially in a long-tailed scenario where the data distribution may vary a lot for each incremental learning step. We address this problem by adding a prediction head to map the current representation space to the past, enabling more efficient knowledge transfer. Replay based methods assume the availability of a memory budget to store a small of number of learned classes data as exemplars to perform knowledge replay during the class-incremental learning. Herding algorithm [59] for exemplar selection based on the class mean vector is firstly proposed in [60], which becomes one of the most popular strategies in existing work [56, 61, 62, 57, 63]. Besides, He et al[42] further improves herding by applying an additional clustering step to select exemplars from each cluster based on herding, which mitigates the issue of intra-class variability. Despite the effectiveness of exemplar replay to maintain the learned knowledge, existing methods assume the training data is balanced and each class has more training data than the memory budget (e.g. 20 exemplars per class). However, the majority of classes in the long-tailed scenario only contain a few training samples, resulting in class-imbalance issues in the exemplar set and may even harm the overall performance when performing the knowledge replay. In this work, we address this issue by constructing a balanced exemplar set by augmenting the tail classes data with the proposed CAM based CutMix, which not only improves the efficiency of knowledge replay to maintain the knowledge but increases the generalization ability on tail classes to achieve better overall performance.
TABLE I: Example of matched food codes in VFN. (NFS denotes Not Further Specified.)
Food Type Food Code Main food description from FNDDS Additional food description
Apple 63101000 Apple, raw apple, NFS
Bagel 51180010 Bagel; flavored bagel; egg bagel; Bagel Thins
Baked potato 71100100 Potato, baked, NFS
Boiled egg 31103010 Egg, whole, boiled or poached soft boiled or hard boiled egg; egg, cooked, no fat or milk added
Chicken wings 24160110 Chicken wing, NFS as to cooking method drummette without sauce or seasoning
Cookies 53201000 Cookie, NFS
Fried rice 58150310 Rice, fried, NFS Chinese rice
Ice cream 13110000 Ice cream, NFS NFS as to flavor
Pancake 55100005 Pancakes, NFS hot cakes; flapjacks
Yogurt 11400000 Yogurt, NFS
III Long-tailed Food Datasets For Diabetes
The most recent work [14] introduced VFN-LT, which is a new long-tailed version of VFN [9] dataset for food recognition where the data distribution exhibits the real world food consumption frequencies [64, 65] for healthy people aged 18 to 65 in U.S.. However, around 34.2 million U.S. individuals (10.5 % of the U.S. population) have diabetes[66], which can cause various health problem such as heart disease, vision loss, and kidney disease. The management of diabetes is significantly influenced by diet, therefore, food recognition for predicting appropriate dietary choices to maintain control of diabetes is essential. Nevertheless, no existing studies have specifically focused on this special population so far. In this work, we aim to fill this gap by introducing two new benchmark long-tailed datasets including VFN-INSULIN and VFN-T2DM, which are constructed based on Viper FoodNet (VFN) datasets [9] and target on food image recognition for dietary assessment among insulin takers and those with type 2 diabetes without taking insulin, respectively.
The original VFN datasets consists of 74 most frequently consumed foods (exclude drink and beverages) selected based on What We Eat In America (WWEIA)  [^1] https://data.nal.usda.gov/dataset/what-we-eat-america-wweia-database. Following the similar process in [14], we first manually match each of the 74 food types in VFN [9] with one 8-digit USDA food codes from the Food and Nutrient Database for Dietary Studies (FNDDS) [^2] https://www.ars.usda.gov/northeast-area/beltsville-md-bhnrc/beltsville-human-nutrition-research-center/food-surveys-research-group/docs/fndds-download-databases/. Each 8-digit USDA food code represents a specific food item reported and consumed in the food supply . Table I shows examples of matched food codes from FNDDS in the VFN dataset. Next, food codes were labeled with their corresponding consumption frequency as determined by Lin et al [64] using the nationally representative dietary data collected through the National Health and Nutrition Examination Survey (NHANES)  [^3] https://www.cdc.gov/nchs/nhanes/index.html from 2009–2016 among U.S adults aged from 20 to 65. The consumption frequency represents how many times a particular food item was reported on one day in the particular age group among the U.S. population. As such, the frequency shows the prominence of the food in comparison with other foods reported as consumed in the population with 774 and 2,758 participants for insulin takers and type 2 diabetes without taking insulin. Finally, we construct VFN-INSULIN and VFN-T2D by reducing the number of training samples for food classes in original VFN based on the matched consumption frequency. Specifically, given the consumption frequency $f_{i}$ for the i-th food class, we calculate the number of training samples by (1)
$s_{i}=n_{i}\times\frac{f_{i}}{f_{max}}$ (1)
where $n_{i}$ is the original number of training samples in VFN for food class $i$ and $s_{i}$ denotes how many training images are kept for this class, which are randomly selected among original training data. $f_{max}$ refers to the maximum matched consumption frequency among the 74 food types in VFN datasets.
Overall, there are 2,056 training images from 74 food classes in VFN-INSULIN with a maximum of 288 training images per class and a minimum of 1 image per class. In addition, there are 2,145 training images in VFN-T2D with a maximum of 288 and a minimum 1 image per class. We observe that the food type Yeast bread, has a very high consumption among the represented adults and dominates the consumption frequency for both insulin takers and type 2 diabetes without taking insulin, resulting in the same imbalance ratio $\rho=\frac{max_{i}\{s_{i}\}}{min_{i}\{s_{i}\}}$ calculated as 288 for both datasets. However, the consumption frequencies of other food types may vary between the two groups. Figure 2 shows the distribution for each food type in VFN-INSULIN and VFN-T2D in descending order based on the number of training samples per class.
Figure 2: The distribution of VFN-INSULIN and VFN-T2D shown in descending order based on number of training samples.
Figure 3: The overview of our proposed framework. The red arrow shows the training procedure where at each incremental learning phase, the input training data contains the new classes images and the exemplars for all seen classes. We leverage the balanced softmax loss [45] $\mathcal{L}_{bs}$ based on the distribution of input training data to learn class-balance representation. We also employed a fixed teacher model obtained from the last incremental step to perform feature-based knowledge distillation $\mathcal{L}_{fkd}$ to address catastrophic forgetting. The blue arrows denote the steps after the training process where we select a small fixed number of $\mathcal{M}$ exemplars per class and store in the memory buffer. Finally, before the next learning phase, we perform the CAM-based exemplar augmentation to construct a balanced exemplar set and combine with the new classes data for the next training phase.
IV Method
In this work, we introduce a novel end-to-end long-tailed continual learning framework for visual food recognition. The overview of our method is shown in Figure 3. To address catastrophic forgetting, we leverage the teacher model learned from the last incremental step and perform feature-based knowledge distillation with an additional prediction head to enable the efficient knowledge transfer. The exemplar set selected based a novel CAM-based data augmentation for tail classes. Finally, we replace the cross-entropy with the balanced softmax loss [45] based on the current training data distribution to learn class-balance visual representation. In this section, we first introduce the preliminaries for continual learning in long-tailed distribution in Section IV-A and then illustrate the detail of each proposed components in Section IV-BIV-C and IV-D, respectively.
IV-A Preliminaries
We focus on continual learning in class-incremental settings where the objective is to learn new classes incrementally and perform classification on all classes seen so far during the inference phase. Specifically, the continual learning in the class-incremental scenario can be formulated as applying an initial model $h_{0}$ to learn a sequence of $N$ tasks denoted as $\mathcal{T}=\{\mathcal{T}^{1},...,\mathcal{T}^{N}\}$ where each task $\mathcal{T}^{i}$ contains $C_{i}$ non-overlapped new classes, which is also known as the incremental step size. During the learning phase of each new task, only the training data $D_{i}=\{\textbf{x}_{i}^{j},y_{i}^{j}\}$ of the current task is available where $\textbf{x}_{i}^{j}$ and $y_{i}^{j}$ denote the $j$-th input image and label, respectively. After each incremental learning step, the updated model $h_{i}$ needs to classify $C_{1:i}$ classes encountered so far. The major challenge of continual learning is catastrophic forgetting [15] where the updated model $h_{i}$ after learning the task $\mathcal{T}^{i}$ forgets the knowledge of previous tasks $\{\mathcal{T}^{1},...,\mathcal{T}^{i-1}\}$, resulting in significant performance degradation to classify $C_{1:i-1}$. In the conventional setup, the training data $D_{i}$ for each task $\mathcal{T}^{i}$ is balanced distributed, containing $|D_{i}|/C_{i}$ samples per class. However, this assumption simplifies the real world complexities especially for food recognition where data is usually long-tailed distributed and exhibits imbalance among food classes. Formally, the training data $D_{i}$ for each task in long-tailed continual learning is a class-imbalanced distributed with each class containing $(0,|D_{i}|)$ training samples. The entire training data $D$ for all the $N$ tasks $\mathcal{T}$ exhibits the long-tail distribution.
IV-A1 Knowledge distillation
Most existing work [55, 62, 60, 61, 56] applies knowledge distillation [18] on output logits to maintain the performance on previously learned classes. Specifically, during the learning step of the task $\mathcal{T}^{i}$, a teacher model $h_{t}=h_{i-1}$ learned from the last task with fixed parameters is employed. The knowledge distillation aims to minimize the difference between the output logits of the current model $L=[o^{1},o^{2},...o^{C_{1:i}}]\in\mathbb{R}^{C_{1:i}\times 1}$ and the outputs of the teacher model $\hat{L}=[\hat{o}^{1},\hat{o}^{2},...\hat{o}^{C_{1:i-1}}]\in\mathbb{R}^{C_{1:i-1}\times 1}$ by
$\mathcal{L}_{kd}=-\sum_{j=1}^{C_{1:i-1}}\hat{L}^{(j)}_{T}log(L^{(j)}_{T})$ (2)
where T is the temperature scalar to learn the hidden knowledge by softening the output distribution as
$\hat{L}_{T}^{(j)}=\frac{\exp{(\hat{o}^{(j)}}/T)}{\sum_{k=1}^{C_{1:i-1}}\exp{(\hat{o}^{(k)}}/T)}$ (3)
Finally, the knowledge distillation is integrated with cross-entropy during the training process by using a hyper-parameter $\alpha$ to learn new classes as well as maintaining the learned knowledge.
$\mathcal{L}=\alpha\mathcal{L}_{kd}+(1-\alpha)\mathcal{L}_{cn}$ (4)
IV-A2 Exemplar replay
As one of the most commonly used strategies to address catastrophic forgetting, the exemplar replay based methods [60, 61, 57] assume the availability of a reasonable memory budget to select a small fixed number of data as exemplars for each seen class and store them in memory buffer (also known as exemplar set). Specifically, after learning each task $\mathcal{T}^{i}$, the lower layers of updated model $h_{i}$ is used to extract feature embeddings for the new classes training data $D_{i}=\{\textbf{x}_{i}^{j},y_{i}^{j}\}$. The Herding algorithm [59] is widely applied to select the most representative data for each new class based on euclidean distance between feature embedding and the class mean vector. Therefore, given a memory budget of $\mathcal{M}$ data per class (also known as memory capacity), a subset of $E_{i}\subseteq D_{i}$ is selected with $|E_{i}|=\mathcal{M}\times C_{i}$ and stored in the memory buffer. Finally, at the beginning of the next new task $\mathcal{T}^{i+1}$, all the exemplars in memory buffer are combined with the new classes training data to construct $E_{i}+D_{i+1}$ for continual learning. In this work, we use Herding as the exemplar selection algorithm while others latest work [42, 67] could also be applied.
IV-B Feature-based Knowledge Distillation
Despite the effectiveness of knowledge distillation in conventional continual learning setup as described in Section IV-A1, it is difficult to apply when new classes exhibit long-tailed distribution since the output logits of the teacher model could be heavily biased towards instance-rich classes due to the severe class-imbalance problem [62, 63]. Directly applying knowledge distillation as in (2) on biased output logits may even harm the overall performance [19]. Therefore, we explore the feature-based knowledge distillation in this work for more efficient knowledge transfer in the long-tailed continual learning scenario. As discussed in [20], one of the challenges when applying feature-based distillation is called the feature space misalignment where the representation of student and teacher models could mismatch in terms of both magnitude and direction. This problem is also relevant in the long-tailed continual learning scenario as the new knowledge is acquired sequentially, which requires the evolution of the continual learning model in the feature space to accommodate the addition of new classes. In this work, we address this challenge by introducing a simple yet effective method as shown in Figure 4. Specifically, instead of directly mimicking the feature from teacher model, we apply an additional predictor $g$ on the head of the continual learning model to map the current representation space to the past in the teacher model. Given an image x, the predictor $g$ takes the feature representation from the current model (i.e. student model) $h_{i}(\textbf{x})$ as input and outputs the mapped feature $g(h_{i}(\textbf{x}))$. Then we distill the knowledge from the teacher model $h_{i-1}$ by
$\mathcal{L}_{fkd}(\textbf{x}))=1-<g(h_{i}(\textbf{x})),h_{i-1}(\textbf{x}))>$ (5)
where <,><,>< , > measures the cosine similarity. By applying the predictor, we provide the student model with more freedom to accommodate the previous learned representation into the current feature space, enabling more efficient knowledge distillation in long-tailed continual learning. The predictor $g$ is be removed after each incremental learning phase. Note that although we apply cosine embedding loss for knowledge distillation, our method can be integrated with other loss functions such as the Mean Squared Error (MSE) loss.
Figure 4: The overview of proposed feature-based knowledge distillation by applying an additional predictor $g$.
IV-C CAM-based Exemplar Augmentation
Existing exemplar replay based methods assume each class should contain at least $\mathcal{M}$ images given $\mathcal{M}$ as the memory budget in Section IV-A2. However, most classes in long-tailed distribution may contain only a few training samples $n<\mathcal{M}$, which imposes two new challenges including (i) inefficiency of knowledge replay due to the insufficient of training samples and (ii) intensification of the class-imbalance issue if we directly combine the stored exemplars with training data from new class due to the imbalanced nature of memory buffer. Therefore, we propose a novel data augmentation method in this work to construct a balanced exemplar set by augmenting the tail classes images to address both aforementioned issues. The overview of proposed data augmentation technique is illustrated in Figure 5. To address the issue of losing semantic information when performing data augmentation [21, 14] as described in Section II-B, we propose to use a class activation map (CAM) [68] to identify the most important region from instance-rare classes images and then preserve the semantic information by performing CutMix [22] to cut and paste the identified region into the images with rich context that are selected based on visual similarity. Specifically, we construct class-balanced memory buffer before each new task $\mathcal{T}^{i+1}$ by augmenting stored images for food classes $C_{t}$ with less than $\mathcal{M}$ exemplars through CutMix in conjunction with images selected from food classes $C_{h}$ containing $\mathcal{M}$ exemplars. Given an input image $\textbf{x}_{t}\in C_{t}$, we first select the most visually similar candidate $\textbf{x}_{h}\in C_{h}$ by comparing the cosine similarity with $h_{i}$ as feature extractor where $\textbf{x}_{h}=\underset{\textbf{x}_{k}\in C_{h}}{\text{argmax}}<h_{i}(\textbf{x}_{t}),h_{i}(\textbf{x}_{k})>$. The lower half of Figure 5 illustrates the procedure to identify the the region to cut and paste into $\textbf{x}_{h}$. Formally, given $\textbf{x}_{t}\in\mathbb{R}^{c\times h\times w}$, the class-activation map $M(\textbf{x}_{t})\in\mathbb{R}^{h\times w}$ is calculated by
$M(\textbf{x}_{t})=\sum_{k}^{d}v_{y_{\textbf{x}_{t}}}^{k}h_{i}(\textbf{x}_{t})$ (6)
where $v_{y_{\textbf{x}_{t}}}\in\mathbb{R}^{d}$ refers to the weight vector in classifier of the current model corresponding to the seen class $y_{\textbf{x}_{t}}\in C_{1:i}$. The value of CAM ranges from $[0,1]$ and a higher value indicates the more discriminative class-specific region. Therefore, we apply a random threshold $\sigma\in(0,1)$ to select the region $M(\textbf{x}_{t})^{T}\in\mathbb{R}^{h\times w}$ where
$M(\textbf{x}_{t})^{T}=\left\{\begin{array}[]{cl}M(\textbf{x}_{t})&{M(\textbf{x}_{t})\geq\sigma}\\ 0&{M(\textbf{x}_{t})<\sigma}\end{array}\right.$ (7)
without losing semantic information of the input image. Finally, we perform CutMix to generate a synthetic image $\tilde{\textbf{x}}_{t}$ by
$\tilde{\textbf{x}}_{t}=(1-S(\textbf{x}_{t})^{T})\odot\textbf{x}_{h}+S(\textbf{x}_{t})^{T}\odot\textbf{x}_{t}$ (8)
where $\odot$ refers to element-wise multiplication and $S(\textbf{x}_{t})^{T}$ denotes the binary mask obtained from $M(\textbf{x}_{t})^{T}$ that 1 indicates the region with $M(\textbf{x}_{t})^{T}>0$. The class label $\tilde{y}_{t}$ of the synthetic image is calculated by the area of the replaced region in $\textbf{x}_{h}$ as
$\tilde{y}_{t}=\frac{1-A_{r}}{A}y_{h}+\frac{A_{r}}{A}y_{t}$ (9)
where $A_{r}$ and $A$ denote the area of the replaced region and the total area of $\textbf{x}_{h}$. $y_{h}$ and $y_{t}$ refer to the original class label for $\textbf{x}_{h}$ and $\textbf{x}_{t}$, respectively. Note that the exemplar augmentation is performed at the beginning of each new task and the augmented images are not stored in the memory buffer. Note that the Grad-CAM[69], which can be regarded as the generalization of CAM [68], could also be applied in our method.
Figure 5: The overview of proposed CAM-based data augmentation technique. The green arrow describes the selection of the most visually similar candidate image and the red arrow illustrates the steps to obtain the most important region of input image to perform CutMix [22].
IV-D Integrated Loss
While the exemplar augmentation mitigates the class-imbalance issue by constructing a balanced memory buffer, the number of available training data between new classes and the stored classes may still vary a lot during the training phase due to the limited memory budget. Existing work [17] addresses this problem by decoupling the training process into two stages to first learn a feature extractor and then fine-tune the classifier using class-balanced sampler. In this work, we propose to use Balanced Softmax (BS) [45] by extending it into a long-tailed continual learning scenario without requiring a decoupled training process. Specifically, during the training phase of the new task $\mathcal{T}^{i}$, a distribution vector $v_{d}\in\mathbb{R}^{C_{1:i}}$ is generated by counting the number of training data of each food class for input images in the current task. Recall $L\in\mathbb{R}^{C_{1:i}}$ is the output logits from current model $h_{i}$, the distribution vector is then used as the prior information when calculating the loss as shown in (10)
$\mathcal{L}_{bs}=\sum_{k=1}^{C_{1:i}}-y^{k}log[\Phi(\bar{v}_{d}^{k}+L^{k})]$ (10)
where $\bar{v}_{d}=v_{d}/sum(v_{d})$ is the normalized distribution vector and $\Phi()$ denotes the $Softmax$ function. Therefore, the larger value in the distribution vector achieves smaller gradients when we compute the cross-entropy using the adjusted logits $v_{t}+L$ and vice versa, which address the class-imbalance issue and enables the end-to-end training pipeline.
The overall training loss function is the weighted sum of feature-based knowledge distillation as described in (5) and the balanced softmax $\mathcal{L}_{bs}$, which can be expressed as
$\mathcal{L}=\mathcal{L}_{bs}+\lambda\mathcal{L}_{fkd}$ (11)
where $\lambda$ is the adaptive ratio to tune the two losses. In this work, as the number of training data $D_{i}$ may vary a lot for each task $\mathcal{T}^{i}$, we propose to calculate $\lambda=\sqrt{|D_{i}|/|D_{1:i}|}$ as the ratio of training data for the current task and the learned tasks. Therefore, the ratio $\lambda$ increases when there are more training data from new classes.
TABLE II: Results on Food101-LT, VFN-LT, VFN-INSULIN and VFN-T2D by comparing with existing continual learning methods in terms of last step accuracy ($A_{L}$) and avergae accuracy $A_{M}$. Best results are marked in bold.
Datasets Food101-LT  VFN-LT    VFN-INSULIN   VFN-T2D
Number of tasks $N=5$ $N=10$ $N=20$ $N=7$ $N=7$ $N=7$
Accuracy (%) $A_{L}$ $A_{M}$ $A_{L}$ $A_{M}$ $A_{L}$ $A_{M}$ $A_{L}$ $A_{M}$ $\quad A_{L}\quad$ $A_{M}$ $A_{L}$ $A_{M}$
LwF [55] 8.62 10.02 5.86 10.10 0.83 3.58 1.02 4.85 4.69 12.35 3.80 7.10
EWC [54] 4.29 5.05 3.70 7.23 0.83 3.96 1.58 6.31 2.97 12.05 1.50 4.47
iCaRL [60] 11.48 12.42 12.46 13.42 11.04 12.77 11.84 18.76 1.91 4.49 10.64 12.36
EEIL [61] 12.88 12.68 10.57 14.63 6.98 12.66 16.87 20.77 14.77 17.36 13.02 12.95
LwM [70] 10.62 10.82 7.22 12.27 2.45 6.82 7.41 12.32 6.24 15.47 4.24 8.13
IL2M [19] 12.55 11.45 10.97 14.21 6.81 12.16 14.87 18.68 14.19 15.24 10.71 11.93
BiC [62] 14.94 16.72 12.39 16.22 10.38 14.36 15.40 20.89 13.53 19.37 15.32 20.70
LUCIR [57] 13.87 16.94 10.17 15.75 6.74 9.32 16.31 23.37 15.35 20.68 13.42 16.38
PODNet [58] 12.04 12.22 10.46 12.07 8.99 12.41 14.75 18.21 14.18 15.06 12.78 15.55
EEIL-2stage [17] 14.16 14.96 13.29 16.38 9.76 14.71 17.61 22.98 13.64 18.60 12.86 14.47
LUCIR-2stage [17] 16.34 18.90 13.03 18.75 10.85 16.53 17.33 24.26 14.49 22.49 12.82 17.85
PODNet-2stage [17] 13.94 17.89 11.12 17.21 10.28 16.28 17.14 25.58 16.04 22.28 14.24 19.71
Ours 17.54 21.83 13.57 19.25 11.78 17.43 20.43 29.33 18.49 27.51 18.54 28.17
V Experiment
In this section, we evaluate our proposed long-tailed continual learning framework for visual food recognition as illustrated in Section IV. Specifically, we first introduce the experimental setup including the split of datasets and implementation detail described in Section V-A and V-B, respectively. Then we compare our method with existing work in Section V-C and conduct an ablation study to show the effectiveness of each individual component in Section V-D. Finally, we discuss potential techniques that can boost the performance for real world food related applications in Section V-E.
V-A Datasets
We conduct experiments on four benchmark long-tailed food image datasets including: (1) Food101-LT [14], (2) VFN-LT [14], (3) VFN-INSULIN and (4) VFN-T2D where the latter two datasets are introduced in Section III.
Food101-LT is the long-tailed version of Food-101 [12] constructed by applying Pareto distribution [71] with the power ratio $\alpha=6$. We randomly partition the 101 food classes into 5, 10, and 20 tasks for continual learning. Therefore, each task comprises 20, 10, and 5 new classes, respectively, with the exception of the first task, which includes one additional class. The test set is kept as balanced with 125 images per class.
VFN-LT, VFN-INSULIN and VFN-T2D are all the long-tailed version of VFN [9] but constructed based on the food consumption frequency of different populations including (i) healthy people (ii) insulin takers and (iii) individuals with type 2 diabetes without taking insulin, respectively. We randomly divide the 74 food classes into $N=7$ tasks with the first task contain 14 new classes and the remaining tasks contain 10 new classes for continual learning. The test set is balanced with 25 images per class. To facilitate an equatable analysis, we use the same testing data in VFN-LT, VFN-INSULIN and VFN-T2D, which is balanced as 25 samples per class with total 1,850 images.
V-B Implementation Detail
Our implementation of neural networks are based on Pytorch framework [72] and we apply the ResNet-18 from scratch as the backbone network for experiments for all datasets. The ResNet implementation follows the setting suggested in [11]. We train 90 epochs for each new task with the learning rate starts from 0.1 and decreased with ratio $1/10$ for every 30 epochs. The batch size is set as 128 and we apply the stochastic gradient descent (SGD) optimizer with weight decay $0.0001$. The memory budget is set as $\mathcal{M}=20$ to store at most 20 images per food class in the memory buffer.
Evaluation protocol: We use Top-1 classification accuracy as the evaluation metric and evaluate the updated model after learning each new task $\mathcal{T}^{i}$ on test data belonging to all classes seen so far $C_{1:i}$. Besides, we report the last step accuracy $A_{L}$ defined as the classification accuracy on the entire test set after learning all $N$ tasks and the average accuracy $A_{M}$ calculated by averaging the accuracy obtained after learning each new task, which shows the overall performance for the entire continual learning procedure. Each experiment is run five times and the average performance is presented.
Existing methods: We compare our proposed framework with existing continual learning methods including LwF [55], EWC [54], iCaRL [60], EEIL [61], LwM [70], IL2M [19], BiC [62], LUCIR [57] and PODNet [58]. Also, we apply the 2-stage framework for long-tailed continual learning in [17] and incorporate it into three existing methods including EEIL-2stage, LUCIR-2stage and PODNet-2stage for comparisons.
Figure 6: Results on VFN-LT, VFN-INSULIN, VFN-T2D and Food101-LT with different number of tasks $N$. Each marker represents the Top-1 classification accuracy evaluated on all classes seen so far after learning each task.
V-C Comparisons With Existing Methods
Table II summarizes the results on Food101-LT, VFN-LT, VFN-INSULIN and VFN-T2D in terms of the last step accuracy $A_{L}$ and the average accuracy $A_{M}$. We observe noticeable improvements for both $A_{L}$ and $A_{M}$ of our proposed method by comparing with existing work especially on VFN-LT, VFN-INSULIN and VFN-T2D, which contains a heavier tail with a larger imbalance ratio as introduced in Section III. For example, we achieve around $5\%$ increase of average accuracy $A_{M}$ and $2\%$ increase of the last step accuracy $A_{L}$ on VFN-INSULIN compared with the 2-stage framework even without requiring the decoupled training process. Furthermore, we obtain the best performance on Food101-LT with different number of tasks $N\in\{5,10,20\}$. We notice both $A_{L}$ and $A_{M}$ accuracy will drop as the total number of tasks $N$ increase since we need to address the catastrophic forgetting to maintain the learned knowledge at each learning phase of new tasks except for the fist task.
Figure 6 shows the results of top-1 classification accuracy evaluated on all the classes seen so far after learning each new task. Our method achieve promising performance at each learning phase of new task. In addition, different from conventional continual learning where the accuracy usually drops monotonously overtime when new tasks arrive due to the forgetting of previous knowledge, we notice that the performance may even increase after learning a new task in the long-tailed scenario. For example on Food101-LT with $N=5$, we observe the increase of accuracy for most methods after learning the third task. This occurs because the number of training samples varies significantly among different tasks in long-tailed continual learning where the model gains better knowledge for tasks with a larger number of training images. However, this imposes new challenges of learning from class-imbalanced data for each new task and also hyper-parameter tuning (e.g. the factor of knowledge distillation term as in Equation 4). In this work, we not only address the catastrophic forgetting by introducing a feature-based knowledge distillation and an effective exemplar augmentation technique but also alleviate the class-imbalance by integrating the balanced softmax with an adaptive ratio to adjust the impact of each loss term as illustrated in Equation 11 and strike a balance of stability-plasticity.
TABLE III: Ablation study on Food101-LT, VFN-LT, VFN-INSULIN and VFN-T2D in terms of average accuracy $A_{M}$.
Food101-LT VFN-LT VFN-INSULIN VFN-T2D
[…]
5.90 8.79 10.55 12.21 11.47 11.93
17.42 15.83 13.96 22.53 21.19 20.73
13.27 12.99 11.64 16.73 15.41 14.93
16.52 14.20 12.02 22.96 21.33 21.05
19.31 17.26 15.49 24.18 22.31 22.19
21.83 19.25 17.43 29.33 27.51 28.17
Figure 7: Examples of augmented food images on VFN-LT using CMO [21], VM-CMO [14], D-Mixup [41] and our proposed CAM-CutMix.
V-D Ablation Study
[…]
We also evaluate our proposed CAM-CutMix by replacing it with existing data augmentation based methods including the CutMix [22] based approaches: (a) the original CutMix used in CMO [21], (b) Visual-Multi CutMix (VM-CMO[14], (c) SnapMix [50] and the Mixup [73] based approach: (d) D-Mixup [41]. We conduct experiments on VFN-LT and Food101-LT with $N=10$ as shown in Table IV. Generally, the CutMix based approaches work better in long-tailed continual learning scenarios than D-Mixup, which is usually applied in multi-label recognition scenarios. In addition, the SnapMix achieve a slightly better performance than CMO and VM-CMO as it also considers the class-activation map (CAM) when generating mixed labels. Our method achieves the best performance as it not only preserves the most important regions based on CAM but also enables seamless CutMix, rather than relying on a randomly generated bounding box. The example augmented food images using VFN-LT are shown in Figure 7. Note that we do not visualize SnapMix [50] as it has the same synthetic image as in CMO [21] but with a different mixed label.
TABLE IV: Ablation study of different data augmentation methods on Food101-LT ($N=10$) and VFN-LT with average accuracy $A_{M}$.
Food101-LT ($N=10$) VFN-LT
CMO [21] 17.28 25.93
VM-CMO [14] 16.47 26.41
SnapMix [50] 18.31 27.62
D-Mixup [41] 15.93 25.14
CAM-CutMix (Ours) 19.25 29.33
V-E Discussions
Despite the performance improvements our proposed framework demonstrates in comparison to existing methods as shown in Table II, the deployment in real-world applications still remains challenging based on the current classification accuracy. Therefore, in this section we discuss potential techniques that could be applied to boost the performance including (1) increaseing the memory buffer capacity to store more exemplar images for knowledge replay and (2) performing transfer learning by pre-training the backbone on large scale image datasets.
V-E1 Memory buffer capacity
As one of the most efficient techniques to address catastrophic forgetting, the performance of knowledge replay greatly relies on the capacity of memory buffer (i.e. and how many exemplar images can be stored). In this part, we evaluate the long-tailed continual learning performance by varying the memory buffer capacity $\mathcal{M}\in\{10,20,30,40,50,100\}$. Figure 8 shows the results in terms of average accuracy $A_{M}$ on Food101-LT ($N=10$) and VFN-LT where we observe consistent performance improvements by increasing the memory capacity. However, the memory buffer capacity is a significant constraint for continual learning in real world application as it requires larger memory storage and also poses challenges related to privacy concerns when storing original images as exemplars. Additionally, we observe a performance bottleneck where increasing memory capacity does not substantially boost performance. This is predominantly due to dual challenges of catastrophic forgetting and class-imbalance problems that arise in the long-tailed continual learning scenario.
Figure 8: Average accuracy ($A_{M}$) on Food101-LT ($N=10$) and VFN-LT by varying the memory buffer capacity $\mathcal{M}\in\{10,20,30,40,50,100\}$.
V-E2 Transfer learning
Applying the deep models pre-trained on large scale image datasets as backbone a network has been a common strategy to enhance performance across a multitude of vision tasks [74, 75, 13]. In this part, we evaluate the efficacy by leveraging a variety of pre-trained models within the context of long-tailed continual learning for visual food recognition. Formally, we consider different network structures with various depth such as ResNet-50 [11], MobileNet [76, 77], EfficientNet [78, 79] Vision Transformers (ViT) [80, 81, 82] and its variants including DeiT [81] and Swin [82] transformers. In addition, we leverage both ImageNet-1K [83] and ImageNet-21K [84] as the pre-training datasets. ImageNet-1K contains 1,000 classes of general objects, which is the subset of full ImageNet-21K that contains 21,841 classes with over 14,197,122 training images. The results on VFN-LT in terms of average accuracy $A_{M}$ is shown in Figure 9. We observe significant performance improvements for over $20\%$ by using pre-trained models on large-scale datasets compared to our results in Table II with a model from scratch. This is mainly attributed to the fact that pre-training enhances the feature extraction capabilities of the backbone network, thereby enabling it to procure the most discriminative features that are essential for downstream tasks. In addition, pre-training on larger scale datasets that contain more images and classes is able to achieve higher accuracy. However, there is trade-off between the computation complexity and the performance where the increase of model parameters would require longer training time and higher computation capability, which may not be practical for specific real world applications with limited resources available. Note that we intentionally refrain from utilizing food datasets for pre-training in this part to prevent potential overlap with any food class in VFN [9]. However, it’s reasonable to anticipate a more substantial performance enhancement if the pre-training were conducted on large-scale food datasets such as Food2K [13].
Figure 9: Average accuracy ($A_{M}$) on VFN-LT by leveraging pre-trained models.
VI Conclusion
In this work, we focus on visual food recognition within the context of long-tailed continual learning, which is closely related to various real world food based applications. We first introduce two benchmark long-tailed food image datasets including VFN-INSULIN and VFN-T2D, which are generated based on VFN to exhibit the real life food consumption frequency for groups with diabetes. In addition, we propose a novel end-to-end framework that is capable of learning new food classes incrementally in long-tailed data distribution without forgetting the learned knowledge. The proposed framework consists of an effective feature-based knowledge distillation structure by leveraging additional prediction head and a novel data augmentation method based on class-activation-map to address both catastrophic forgetting and class-imbalance issues simultaneously. Our method shows improved performance on all datasets, Food101-LT, VFN-LT, VFN-INSULIN and VFN-T2D, comparing with existing work. Finally, we conduct an ablation study to evaluate the effectiveness of each individual component and discuss the trade-off of potential techniques to achieve higher accuracy including the memory buffer capacity and transfer learning. Our future work includes designing an exemplar-free long-tailed continual learning framework. If successful, this could address the current challenges associated with large memory buffer and potential privacy concerns associated with storing food images.